OpenStudy (anonymous):

The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (-3, 2).

OpenStudy (shamil98):

slope of the line perpendicular is the negative reciprocal what do you get as the slope then?

OpenStudy (kc_kennylau):

These questions may help you think: (a) What is the slope of LM? (b) What is the slope of the line perpendicular to the line LM? (c) You now have the slope and a point, can you express the equation of the perpendicular line in point-slope form? (d) Can you convert that to slope-intercept form? Wish you all the best :D

OpenStudy (anonymous):

Select one: a. y = -5x + 13 b. y= -1/5x + 7/5 c. y= 1/5x -7/5 d. y= 5x-15

OpenStudy (isaiah.feynman):

Solve it and select the option yourself sir.